# A hidden Markov model for statistical arbitrage in international crude oil futures markets

### v2

@@ 3.2. Performance measurement
As pointed out in [Gatev et al., 2006], computing returns for trading strategies based on the concept of pairs trading is non-trivial due to several factors that complicate the calculation and interpretation of performance metrics. First of all, there are days in which a trading position is open and days in which there is no open position. Moreover, a successfully executed trade delivers a positive cashflow when opened and a non-negative cashflow when closed, thus preventing the computation of a standard linear return. To overcome these issues, it is common to rely on the daily mark-to-market profit & loss indicator.
Let us assume that the strategy is applied over a time period spanning $n$ days. Following [Burgess, 1999], for each day $t=1,\ldots,n$ we compute a daily return $r_{t}$ in the following way:
  • if there is no open position at $t$, then $r_{t}:=0$;
  • if there is an open position at $t$, then
    (3.1) $r_{t}:=S_{t}\frac{S_{t}-S_{t-1}}{V_{t}^{S}}-c\left|S_{t}-S_{t-1}\right|,$
    where $V_{t}^{S}:=|\lambda^{0}|+\sum_{i=B,S,W}|\lambda^{i}|F_{t}^{i}$ represents the total market exposure of the portfolio and $c$ accounts for proportional transaction costs, as explained below.
@@
which can be interpreted as the return over a time period spanning $n$ days. As in our empirical analysis we will consider time periods of different lengths, we shall always report the returns in (3.2) in annual terms: following the usual convention, this amounts to multiply $R$ by $250/n$.
When evaluating the performance of statistical arbitrage strategies, transaction costs can play a significant role, as discussed in [Do and Faff, 2012]. Indeed, these strategies typically involve frequent trading and possibly considerable margins to be posted. Following a common approach in the literature, we take into account these market frictions by introducing in (3.1) the parameter $c$ representing proportional transaction costs. This parameter is however difficult to estimate. Considering the literature on pairs trading on USD-denominated crude oil futures, we found a minimum value of $c=20$bps in [Alizadeh and Nomikos, 2008] and a maximum value of $c=60$bps in [Do and Faff, 2012]. In the present work, we consider trading strategies that include the Shanghai futures, which is denominated in CNY and therefore incurs into currency exchange costs. When including the Shanghai futures in our statistical arbitrage strategies, we assume $c=80$bps. This generates a significant amount of transaction costs, which will therefore lead to rather conservative performance measures for our trading strategies.
@@
Unlike pure arbitrage opportunities, statistical arbitrage strategies are not exempt of financial risk. For this reason, we shall also consider the Sharpe ratio, a widely used risk-adjusted performance measure:
(3.3) $SR:=\frac{\frac{1}{n}\sum_{t=1}^{n}r_{t}}{\sqrt{\frac{1}{n}\sum_{t=1}^{n}\left(r_{t}-\frac{1}{n}\sum_{v=1}^{n}r_{v}\right)^{2}}}.$
@@
For the sake of comparability, we shall always express the Sharpe Ratio of a given trading strategy in annual terms. In Section 4.2.1, we will perform an additional analysis by also computing the Value-at-Risk of the different strategies introduced in Section 3.1.
@@ 4.1.1. Data description
We consider daily [^5] Statistical arbitrage opportunities arise also intraday. See for instance [Marshall et al., 2013] for an analysis of this kind of opportunities in the ETFs market. and weekly futures prices at one month maturity for the Brent, the WTI and the Shanghai futures, the latter converted from CNY to USD. While the time series for the Brent and the WTI start in the eighties, the Shanghai has been trading only since March 2018. Therefore, our dataset spans from $t_{0}=$ 03/26/2018 to $T=$ 06/30/2023, including 1374 daily observations and 274 weekly ones. The weekly futures prices are displayed in Figure 1.
@@ 4.1.2. Unit roots, VAR and cointegration analysis
ADF 0.9434 0.8507 0.9422
PP 0.8790 0.8386 0.8711
bP 0.9548 0.9402 0.9416
Table 1. P-values of the Augmented Dickey-Fuller (ADF), the Phillips-Perron (PP) and the breakpoint Perron (bP) tests. The number of lag length for the ADF and the bP is chosen according to the Schwarz IC.
@@
$\left[\begin{array}[]{c}\underset{}{F_{t}^{B}}\\ \underset{}{F_{t}^{S}}\\ \underset{}{F_{t}^{W}}\end{array}\right]=\left[\begin{array}[]{c}\underset{(1.095)}{1.546}\\ \underset{(0.883)}{1.111}\\ \underset{(1.093)}{1.353}\end{array}\right]+\left[\begin{array}[]{ccc}\underset{(0.249)}{0.614^{**}}&\underset{(0.125)}{0.088}&\underset{(0.234)}{{0.147}}\\ \underset{(0.201)}{0.510^{**}}&\underset{(0.101)}{0.555^{***}}&\underset{(0.189)}{{-0.100}}\\ \underset{(0.249)}{-0.221}&\underset{(0.125)}{0.111}&\underset{(0.233)}{{0.968^{***}}}\end{array}\right]\left[\begin{array}[]{c}\underset{}{F_{t-1}^{B}}\\ \underset{}{F_{t-1}^{S}}\\ \underset{}{F_{t-1}^{W}}\end{array}\right]$
$\qquad+\left[\begin{array}[]{ccc}\underset{(0.258)}{0.266}&\underset{(0.113)}{-0.157}&\underset{(0.238)}{0.033}\\ \underset{(0.208)}{-0.221}&\underset{(0.091)}{0.157^{*}}&\underset{(0.192)}{0.078}\\ \underset{(0.257)}{0.242}&\underset{(0.112)}{-0.214^{*}}&\underset{(0.237)}{0.100}\end{array}\right]\left[\begin{array}[]{c}\underset{}{F_{t-2}^{B}}\\ \underset{}{F_{t-2}^{S}}\\ \underset{}{F_{t-2}^{W}}\end{array}\right]+\left[\begin{array}[]{c}\underset{}{u_{t}^{B}}\\ \underset{}{u_{t}^{S}}\\ \underset{}{u_{t}^{W}}\end{array}\right]$
where ∗∗∗, ∗∗ and represent, respectively, a statistical significance at the 1%, 5% and 10% level. Inspecting the roots of the characteristic equation, this VAR(2) appears to satisfy the stability condition. We can now perform the cointegration test. We adopt the trace test of [Johansen, 1988] and allow for a constant in the cointegrating relationship.
r h Trace test stat. Crit. val. pValue
0 1 36.107 35.193 0.039
1 0 10.648 20.262 0.603
2 0 1.250 9.164 0.915
SBIC($r=1$): 12.9533
@@
Since we cannot reject that $\mathrm{rank}\left[\mathbf{\Pi}_{0}\right]=1$, the matrix $\mathbf{\Pi}_{0}$ can be decomposed as $\mathbf{\Pi}_{0}=\boldsymbol{\alpha}\boldsymbol{\beta}^{T}$ with $\boldsymbol{\beta}\in\mathbb{R}^{3}$. Rewriting $\boldsymbol{\pi}_{0}$ as $\boldsymbol{\pi}_{0}=\boldsymbol{\alpha}c_{0}$, with $c_{0}\in\mathbb{R}$, we can derive the expression of the resulting cointegration relationship, $c_{0}+\boldsymbol{\beta}^{\top}{}\mathbf{F}_{t}$, which is our spread process $S$, and the adjustment coefficients $\boldsymbol{\alpha}$ that measure the speed of convergence of each series to the long-run relationship. Therefore, in our test sample the spread process results to be given by
@@
The estimated adjustment coefficients of the VECM are $\boldsymbol{\alpha}=\left[0.0640\quad 0.4226\quad 0.0684\right]^{\top}$. It is interesting to observe that the Shanghai futures price realigns to the long-run equilibrium much faster than the other two futures prices. As a high speed of adjustment in one or more of the portfolio constituents is a desirable property when constructing statistical arbitrage strategies, this provides a preliminary indication of the potential profitability of including the Shanghai futures in the context of crude oil futures portfolios.
@@ 4.1.3. Filtering the spread
We now estimate the parameters of the OU-HMM for $S$. We first determine the most likely number $N$ of the states of the latent Markov chain X; we then compute daily filtered estimates of X; finally, we estimate the model parameters $\boldsymbol{\gamma}$, $\boldsymbol{\alpha}$, $\boldsymbol{\eta}$ in (2.3) and the transition matrix $\boldsymbol{\Pi}$ by means of the filter-based EM algorithm described in Section 2.2. We implement the filtering and parameter estimation procedure on daily data. The choice of daily data is motivatied by two reasons: first, we want to rely on a large dataset for the estimation of the parameters of the spread process, in order to capture its short-term variations; second, we want to possibly implement trading strategies on a daily basis rather than on a weekly basis.
The daily spread process is determined by using the weights reported in (4.1). We estimate the whole model for $N\in\{1,2,3\}$ and compute the information criteria and the mean square error at the end of the training sample. Although most of the times there is complete agreement among the different information criteria on the optimal value of $N$, we follow the Schwarz IC. According to this criterion, we find that in our training sample the best model entails $N=2$ states of the Markov chain X.
@@
Figure 4. Recursively updated estimates of $\gamma_{i}$, $\alpha_{i}$, $\eta_{i}$ and $\pi_{ii}$, $i=1,2$, for the $N=2$ OU-HMM.
Due to the well-known issue of slow convergence of the EM algorithm, the estimates turn out to be unstable at the beginning of the training sample. Afterwards, they stabilize and only major changes in the underlying observations have an impact on the estimates.
state 1 state 2 state 1 state 2
$\gamma$ 0.1632 0.0869 $\beta$ 0.8490 0.3233
$\alpha$ 0.8077 0.7313 $a$ 53.3830 78.2361
$\eta$ 1.4312 1.4832 $\xi$ 25.0833 27.2027
$\pi_{11}$ 0.7138 $\pi_{22}$ 0.6640
Table 3. Point estimates of the functions of the OU-HMM (both in discrete and continuous time) at the end of the training sample.
For the sake of completeness, the parameter estimates at the end of the training sample are reported in Table 3. As we can see, the main feature that differentiates the two regimes is the value of $\gamma$: the two regimes display almost the same level of persistence $\alpha$, the same level of volatility $\eta$ but a different mean level $\gamma$. This difference turns out to be even more striking in the other training samples considered below and most of the times we find $\gamma_{1}>0$ and $\gamma_{2}<0$.
@@
Figure 5. One-step ahead forecasts and actual values of $y$ across the test sample.

### v3

@@ 3.2. Performance measurement
As pointed out in [Gatev et al., 2006], computing returns for trading strategies based on the concept of pairs trading is non-trivial due to several factors that complicate the calculation and interpretation of performance metrics. First of all, there are days in which a trading position is open and days in which there is no open position. Moreover, a successfully executed trade delivers a positive cashflow when opened and a non-negative cashflow when closed, thus preventing the computation of a standard return. To overcome these issues, it is common to rely on the daily mark-to-market profit & loss indicator.
Let us assume that the strategy is applied over a time period spanning $n$ days. Following the original methodology proposed by [Gatev et al., 2006], which is widely used in the context of pairs trading strategies (see, e.g., [Cummins and Bucca, 2012] and [Hain et al., 2018]), we compute the daily mark-to-market return $r_{t}$ for each day $t=1,\ldots,n$ as
(3.1) $r_{t}=\sum_{k=1}^{K}\left(I_{t}^{k}\cdot\frac{F_{t}^{k}-F_{t-1}^{k}}{F_{t-1}^{k}}\cdot\frac{F_{t-1}^{k}|\lambda^{k}|}{|\lambda^{0}|+\sum_{j=1}^{K}F_{t-1}^{j}|\lambda^{j}|}-|\Delta I_{t}^{k}|\cdot\frac{c_{k}F_{t}^{k}|\lambda^{k}|}{|\lambda^{0}|+\sum_{j=1}^{K}F_{t-1}^{j}|\lambda^{j}|}\right)$
where:
  • $K$ is the number of futures contracts considered in the trading strategy, which is three in our benchmark case;
  • $I_{t}^{k}$ is a three-state indicator variable that takes the value $1$, 1-1, or 0, corresponding to a long, short, or neutral (no position) state, respectively, for futures contract $k$. $\Delta I_{t}^{k}=I_{t}^{k}-I_{t-1}^{k}$, which is nonzero only when a position is opened or closed on day $t$;
  • $c_{k}$ represents the transaction cost parameter associated with trading futures contract $k$, applied when a position is opened or closed.
The first term in (3.2) represents the daily mark-to-market profit or loss generated by the futures positions, weighted by their relative contribution to the total gross exposure. The second term accounts for transaction costs incurred when positions are opened or closed, proportional to the current market value of the traded contracts and independent of trading direction. By construction, if there is no open position on day $t$, then $r_{t}=0$. The daily returns defined in (3.2) represent excess returns per unit of gross notional exposure, as in [Hain et al., 2018]. In particular, the normalization ensures comparability across different strategies. Because trading positions are initiated only when the spread deviates sufficiently from its equilibrium value, the opening of a position generates an immediate positive cash inflow, while closing positions produces no negative cash outflow (apart from transaction costs). As a consequence, the daily returns defined in (3.2) can be interpreted as excess returns.
@@
To facilitate comparison across samples of different lengths, we shall always report the returns in (3.2) in annual terms: following the usual convention, this amounts to multiply $R$ by $250/n$.
Daily returns computed as above typically alternate between periods in which a position is active and periods in which no position is held. As a consequence, daily returns may exhibit substantial short-term variability and are of limited standalone relevance for an investor primarily concerned with the profitability over complete trading episodes. For this reason, we introduce an additional trade-level performance measure based on individual trading windows, defined as the time interval between the opening and closing of a position. Formally, let $I^{k}_{t}\in\{-1,0,1\}$ denote the position indicator for futures contract $k$ at time $t$, as defined above, and define the aggregate position indicator
$I_{t}:=\max_{k=1,\dots,K}|I^{k}_{t}|,$
which equals one whenever at least one futures position is open and zero otherwise. Let $\{(\sigma_{j},\tau_{j})\}_{j=1}^{N_{\mathrm{TW}}}$ denote the resulting sequence of non-overlapping trading windows, where $N_{\mathrm{TW}}$ denotes the total number of completed trading windows over the entire sample and
$\sigma_{j}:=\min\{t>\tau_{j-1}:I_{t}=1,\ I_{t-1}=0\}\quad\text{ and }\quad\tau_{j}:=\min\{t>\sigma_{j}:I_{t}=0\},$
with $\tau_{0}:=0$. For each trade $j=1,\ldots,N_{\mathrm{TW}}$, the trading-window return is defined as
(3.3) $R^{\mathrm{TW}}_{j}:=\prod_{t=\sigma_{j}+1}^{\tau_{j}}\bigl(1+r_{t}\bigr)-1,$
where $r_{t}$ denotes the daily return defined in (3.1). For brevity of notation, we denote by $R^{\mathrm{TW}}$ the collection of trading-window returns over the entire sample. Note that, since $r_{t}=0$ whenever no position is open, $R^{\mathrm{TW}}_{j}$ coincides with the compounded return generated over the holding period of trade $j$. Trading-window returns thus provide a direct measure of per-trade profitability, complementary to daily and aggregate return measures.
When evaluating the performance of statistical arbitrage strategies, which typically entail frequent trading, transaction costs can play a significant role and substantially affect net profitability, as discussed in [Do and Faff, 2012]. Following [Hain et al., 2018], we estimate each contract-specific transaction cost parameter $c_{i}$ as the average effective daily half bid-ask spread over the full sample period. The resulting transaction cost estimates are reported in Table 1. As expected, the highest transaction cost is associated with the Shanghai contract, which corresponds to the youngest and least mature market among those considered and may therefore be subject to relatively higher trading frictions.
Futures contract Transaction costs parameter $c_{i}$ (bps)
Brent 5.80
Shanghai 53.71
WTI 20.24
Dubai 1.52
Table 1. Estimates of the transaction cost parameters $c_{i}$ for the four futures contracts considered in our statistical arbitrage strategies.
Overall, the estimated transaction cost parameters are in line with values documented in the literature. In particular, studies on pairs trading involving USD-denominated crude oil futures report transaction costs ranging from approximately $c=20$ bps in [Alizadeh and Nomikos, 2008] to about $c=60$ bps in [Do and Faff, 2012].
@@
Unlike pure arbitrage opportunities, statistical arbitrage strategies are not exempt of financial risk. For this reason, we also consider the Sharpe ratio, a widely used risk-adjusted performance measure
(3.4) $SR:=\frac{\frac{1}{n}\sum_{t=1}^{n}r_{t}}{\sqrt{\frac{1}{n}\sum_{t=1}^{n}\left(r_{t}-\frac{1}{n}\sum_{v=1}^{n}r_{v}\right)^{2}}},$
@@
where we do not subtract the prevailing risk-free rate in the numerator, as our returns $r_{t}$ represent excess returns. For the sake of comparability, we shall always express the Sharpe Ratio of a given trading strategy in annual terms. In Section 4.2.1, we will perform an additional analysis by also computing the Value-at-Risk, the Expected Shortfall and the Maximum Drawdown of the different strategies introduced in Section 3.1. Moreover, to address the limitation that the Sharpe Ratio captures only the first two moments of the return distribution, we also examine the full empirical distribution of returns generated by our strategies.
@@ 4.1.1. Data description
We consider daily [^5] Statistical arbitrage opportunities arise also intraday. See for instance [Marshall et al., 2013] for an analysis of this kind of opportunities in the ETFs market. and weekly futures prices for the Brent, the WTI and the Shanghai futures, the latter converted from CNY to USD. Specifically, we focus on continuous price series constructed from futures contracts with the nearest monthly maturity, identified by the following Datastream mnemonics: LLCC.01 for Brent, NCLC.01 for WTI, SICC.01 for Shanghai, and OLDUB1M for Dubai crude oil futures, which is introduced in the analysis in Subsection 4.3. By construction, these rolled series may exhibit a non-negligible roll yield, arising from the fact that the price of an expiring contract can differ from that of the next-nearest maturity. However, as discussed in Section 2, our statistical arbitrage strategies are based on the cointegrating relationships among these futures. Since all the futures contracts are written on essentially the same underlying asset (crude oil), the estimated cointegrating coefficients tend to sum to a value close to zero. This property naturally mitigates the impact of roll yield, as the roll yields across different contracts usually move together and, when combined linearly with cointegrating coefficients that approximately sum to zero, their effects largely offset one another.
While the time series for the Brent and the WTI start in the eighties, the Shanghai has been trading only since March 2018. Therefore, our dataset spans from $t_{0}=$ 03/26/2018 to $T=$ 06/30/2023, including 1373 daily observations and 273 weekly ones. The weekly futures prices are shown in Figure 1, where the sharp decline in crude oil prices following the COVID-19 crisis in early 2020 as well as the pronounced peak associated with the outbreak of the Russia-Ukraine conflict in 2022 and the subsequent period of increased volatility are clearly visible.
@@ 4.1.2. Unit roots, VAR and cointegration analysis
ADF 0.9434 0.8535 0.9423
$P$ 0.8788 0.8413 0.8714
bP 0.9548 0.9471 0.9425
Table 2. $p$-values of the Augmented Dickey-Fuller (ADF), the Phillips-Perron ($P$) and the breakpoint Perron (bP) tests. The number of lag length for the ADF and the bP is chosen according to the Schwarz IC.
@@
$\left[\begin{array}[]{c}F_{t}^{B}\\ F_{t}^{S}\\ F_{t}^{W}\end{array}\right]$ $=\left[\begin{array}[]{c}\underset{(1.095)}{1.546}\\ \underset{(0.883)}{1.111}\\ \underset{(1.093)}{1.354}\end{array}\right]+\left[\begin{array}[]{ccc}\underset{(0.249)}{0.614^{**}}&\underset{(0.125)}{0.088}&\underset{(0.234)}{0.147}\\ \underset{(0.201)}{0.510^{**}}&\underset{(0.101)}{0.555^{***}}&\underset{(0.189)}{-0.100}\\ \underset{(0.249)}{-0.221}&\underset{(0.125)}{0.111}&\underset{(0.233)}{0.968^{***}}\end{array}\right]\left[\begin{array}[]{c}F_{t-1}^{B}\\ F_{t-1}^{S}\\ F_{t-1}^{W}\end{array}\right]$
$\quad+\left[\begin{array}[]{ccc}\underset{(0.258)}{0.266}&\underset{(0.113)}{-0.157}&\underset{(0.238)}{0.034}\\ \underset{(0.208)}{-0.221}&\underset{(0.091)}{0.158^{*}}&\underset{(0.192)}{0.077}\\ \underset{(0.257)}{0.242}&\underset{(0.112)}{-0.214^{*}}&\underset{(0.237)}{0.100}\end{array}\right]\left[\begin{array}[]{c}F_{t-2}^{B}\\ F_{t-2}^{S}\\ F_{t-2}^{W}\end{array}\right]+\left[\begin{array}[]{c}u_{t}^{B}\\ u_{t}^{S}\\ u_{t}^{W}\end{array}\right]$
where ∗∗∗, ∗∗ and represent, respectively, a statistical significance at the 1%, 5% and 10% level. Inspecting the roots of the characteristic equation, this VAR(2) specification appears to satisfy the stability condition.
We can now perform the cointegration test. We adopt the trace test of [Johansen, 1988] and allow for a constant in the cointegrating relationship.
r h Trace test stat. Crit. val. $p$-value
0 1 35.5106 29.7976 0.0099
1 0 10.0981 15.4948 0.3074
2 0 0.8131 3.8415 0.5153
@@
Since we cannot reject that $\mathrm{rank}\left[\mathbf{\Pi}_{0}\right]=1$, the matrix $\mathbf{\Pi}_{0}$ can be decomposed as $\mathbf{\Pi}_{0}=\boldsymbol{\alpha}\boldsymbol{\beta}^{T}$ with $\boldsymbol{\beta}\in\mathbb{R}^{3}$. Rewriting $\boldsymbol{\pi}_{0}$ as $\boldsymbol{\pi}_{0}=\boldsymbol{\alpha}c_{0}+\mathbf{c}_{1}$, with $c_{0}\in\mathbb{R}$ and $\mathbf{c}_{1}\in\mathbb{R}^{3}$, we can derive the expression of the resulting cointegration relationship, $c_{0}+\boldsymbol{\beta}^{\top}{}\mathbf{F}_{t}$, which is our spread process $S$, and the adjustment coefficients $\boldsymbol{\alpha}$ that measure the speed of convergence of each series to the long-run relationship. The maximum likelihood estimates of the full VECM, together with their standard errors and statistical significance, indicated as before by ∗∗∗, ∗∗, and , are
(4.10) $\left[\begin{array}[]{c}\Delta F_{t}^{B}\\ \Delta F_{t}^{S}\\ \Delta F_{t}^{W}\end{array}\right]$ $=\left[\begin{array}[]{c}\underset{(0.271)}{0.200}\\ \underset{(0.217)}{0.131}\\ \underset{(0.269)}{0.199}\end{array}\right]+\left[\begin{array}[]{c}\underset{(0.151)}{0.059}\\ \underset{(0.121)}{0.419^{***}}\\ \underset{(0.149)}{0.063}\end{array}\right]\Big(F_{t-1}^{B}-0.698\,F_{t-1}^{S}-0.340\,F_{t-1}^{W}+0.432\Big)$
(4.17) $\quad+\left[\begin{array}[]{ccc}\underset{(0.267)}{-0.362}&\underset{(0.142)}{0.088}&\underset{(0.235)}{0.140}\\ \underset{(0.214)}{0.109}&\underset{(0.114)}{-0.162}&\underset{(0.188)}{0.039}\\ \underset{(0.265)}{-0.220}&\underset{(0.141)}{0.093}&\underset{(0.234)}{0.007}\end{array}\right]\left[\begin{array}[]{c}\Delta F_{t-1}^{B}\\ \Delta F_{t-1}^{S}\\ \Delta F_{t-1}^{W}\end{array}\right]$
(4.27) $\quad+\left[\begin{array}[]{ccc}\underset{(0.263)}{-0.132}&\underset{(0.119)}{-0.073}&\underset{(0.239)}{0.267}\\ \underset{(0.211)}{-0.103}&\underset{(0.095)}{-0.133}&\underset{(0.191)}{0.144}\\ \underset{(0.261)}{0.111}&\underset{(0.118)}{-0.102}&\underset{(0.237)}{0.061}\end{array}\right]\left[\begin{array}[]{c}\Delta F_{t-2}^{B}\\ \Delta F_{t-2}^{S}\\ \Delta F_{t-2}^{W}\end{array}\right]+\left[\begin{array}[]{c}\varepsilon_{t}^{B}\\ \varepsilon_{t}^{S}\\ \varepsilon_{t}^{W}\end{array}\right].$
Having framed the system within a VECM framework, it is possible to investigate both short-run and long-run Granger causality, as well as their direction, among the variables. The estimates reported in (4.17) provide limited evidence of short-run causality, since the coefficients on the lagged first-difference terms are generally not statistically significant at conventional levels. In contrast, the adjustment coefficient in the equation for $\Delta F_{t}^{S}$ is positive and statistically significant at the 1% level, while the corresponding coefficients in the equations for $\Delta F_{t}^{B}$ and $\Delta F_{t}^{W}$ are not statistically different from zero. This pattern indicates that deviations from the long-run equilibrium are primarily corrected through adjustments in $F_{t}^{S}$, implying a unidirectional long-run causal relationship running from $F_{t}^{B}$ and $F_{t}^{W}$ to $F_{t}^{S}$. From a trading perspective, the fact that the Shanghai futures price adjusts to restore the long-run relationship suggests that it plays a stabilizing role within the system, making it a promising candidate for pairs trading strategies based on mean reversion.
Given the estimates in (4.17), in our test sample the spread process results to be given by
@@ 4.1.3. Filtering the spread
We now estimate the parameters of the AR-HMM for $S$. We first determine the most likely number $N$ of the states of the latent Markov chain X; we then compute daily filtered estimates of X; finally, we estimate the model parameters $\boldsymbol{\gamma}$, $\boldsymbol{\alpha}$, $\boldsymbol{\eta}$ in (2.2) and the transition matrix $\boldsymbol{\Pi}$ by means of the filter-based EM algorithm described in Section 2.2. We implement the filtering and parameter estimation procedure on daily data. The choice of daily data is motivatied by two reasons: first, we want to rely on a large dataset for the estimation of the parameters of the spread process, in order to capture its short-term variations; second, we want to possibly implement trading strategies on a daily basis rather than on a weekly basis.
The daily spread process is determined by using the weights reported in (4.28). We estimate the model for $N\in\{1,2,3\}$. Since our trading strategy is based on the forecasts generated by the model, we compute the mean forecast error over the trading sample and select the value of $N$ that minimizes it. According to this criterion, the preferred specification in the training sample features $N=2$ states for the Markov chain $\mathbf{X}$.
@@
Figure 4. Recursively updated estimates of $\gamma_{i}$, $\alpha_{i}$, $\eta_{i}$ and $\pi_{ii}$, $i=1,2$, for the $N=2$ AR-HMM.
Due to the well-known issue of slow convergence of the EM algorithm, the estimates turn out to be unstable at the beginning of the training sample. Afterwards, they stabilize and only major changes in the underlying observations have an impact on the estimates. For example, we observe an instability in the filtered parameter estimates around the onset of the war in Ukraine, coinciding with a period of heightened market volatility.
For the sake of reference, row “baseline” of Table 4 reports the parameter estimates at the end of the training sample. The results indicate that both regimes are highly persistent, with the second regime being particularly so, as reflected by a transition probability $\pi_{22}$ close to one. The two regimes exhibit comparable mean-reversion speeds, as measured by the autoregressive coefficients $\alpha$, while differing markedly in their long-run mean and volatility parameters. In particular, regime 1 is characterized by higher values of the long-run mean level $\gamma$ and volatility $\eta$, whereas regime 2 exhibits markedly lower values for both parameters.
@@
Figure 5. One-step ahead forecasts and actual values of $S$ across the test sample.
@@
State 1 State 2
$\hat{\pi}_{11}$ $\hat{\pi}_{22}$ $\hat{\gamma}_{1}$ $\hat{\alpha}_{1}$ $\hat{\eta}_{1}$ $\hat{\gamma}_{2}$ $\hat{\alpha}_{2}$ $\hat{\eta}_{2}$
baseline 0.8588 0.9732 0.4705 0.6214 1.8343 0.0855 0.8164 1.2187
mean 0.9255 0.9627 0.4504 0.6426 1.5606 0.0600 0.8049 1.2777
std. dev. 0.0583 0.0559 0.0948 0.1319 0.4821 0.0587 0.0405 0.0951
$t$-ratio 15.87 17.21 4.75 4.87 3.24 1.02 19.85 13.44
Table 4. Point estimates of the AR-HMM parameters at the end of the training sample under different initializations. The baseline case refers to the initialization procedure described in part (1) of Remark 2.2. The “mean” and “std. dev.” rows summarize parameter estimates from 100 runs of the filtering algorithm with random initializations, as described in Remark 4.1.
Remark 4.1 (Robustness of the estimates to filtering initialization).
To assess the stability of the parameter estimates with respect to the initialization of the filtering algorithm, we consider the two-state HMM specification and repeat the filter-based EM algorithm under randomized initial conditions. This procedure also serves to address potential local-maxima issues inherent in EM estimation (see, e.g., [McLachlan and Krishnan, 2008, Chapter 3]). We perform 100 independent runs of the algorithm. The transition probabilities $\pi_{11}$ and $\pi_{22}$ are independently drawn from the interval $(0.5,1)$ in order to enforce a minimum degree of regime persistence in the hidden Markov chain. The initial guesses for the regime-dependent AR parameters are anchored to the preliminary OLS estimates obtained as in part (1) of Remark 2.2. For each parameter, lower and upper bounds are drawn uniformly from $(0.5,0.95)$ and $(1.05,1.5)$ times the corresponding OLS estimate, respectively. States are consistently labeled so that State 1 is associated with the larger value of $\gamma$. The results reported in Table 4 indicate a remarkable stability in the estimated transition probabilities and limited variability in most regime-dependent parameters [^6] In Table 4, the $t$-ratio (mean divided by standard deviation) provides a quantification of estimation uncertainty.. For the majority of parameters, the mean estimate across initializations substantially exceeds the corresponding standard deviation, indicating limited sensitivity to initialization. An exception is the long-run mean in the second regime, whose estimate is close to zero. As documented in Remark 4.3, these mild instabilities have negligible impact on the performance of the statistical arbitrage strategies, which are primarily driven by opening signals and are therefore less sensitive to small deviations in the model parameters. This indicates that the economically relevant outcomes of the model are robust to initialization and not driven by convergence to a specific local optimum.